Calculate the normality sulphuric acid solution if mole fraction of water in sulphuric acid solution is 0.85

Grade:12

4 Answers

Arun
5 years ago

We have mole fraction of solute + solvent = 1

Here, Mole fraction of water is 0.85 so mole fraction of sulphuric acid will be (1-0.85) = 0.15

Hence each mole of the solution, 0.15 mole of sulphuric acid is dissolved .

So mass of solvent (water) is ,

For 1 mole of water (H2O) = 18 gm

So, 0.85 mole of water = 18 x 0.85 = 15.3 gm =0.0153 Kg

Thus, Molality (m) = no. of moles of solute (Sulphuric acid)/ mass of solvent in Kg

Hence, Molality of the solution (m) = 0.15 / 0.0153 = 9.8 m (nearly equal to 10)

 

Regards

Arun

Gitanjali Rout
5 years ago
 

We have mole fraction of solute + solvent = 1

Here, Mole fraction of water is 0.85 so mole fraction of sulphuric acid will be (1-0.85) = 0.15

Hence each mole of the solution, 0.15 mole of sulphuric acid is dissolved .

So mass of solvent (water) is ,

For 1 mole of water (H2O) = 18 gm

So, 0.85 mole of water = 18 x 0.85 = 15.3 gm =0.0153 Kg

Thus, Molality (m) = no. of moles of solute (Sulphuric acid)/ mass of solvent in Kg

Hence, Molality of the solution (m) = 0.15 / 0.0153 = 9.8 m (nearly equal to 10)

Gitanjali Rout
5 years ago

Hence each mole of the solution, 0.15 mole of sulphuric acid is dissolved .

So mass of solvent (water) is ,

For 1 mole of water (H2O) = 18 gm

So, 0.85 mole of water = 18 x 0.85 = 15.3 gm =0.0153 Kg

Thus, Molality (m) = no. of moles of solute (Sulphuric acid)/ mass of solvent in Kg

Hence, Molality of the solution (m) = 0.15 / 0.0153 = 9.8 m (nearly equal to 10)

Abhiraj Pathak
5 years ago
Normality = gm equivalent of solute/Vol of solution (L)Considering 100g of water, i.e. 100/18 = nb =5.55 moles of water,xb = 0.85 = nb/(nb+na) On solving we get na = 0.979, i.e. 95.98 g of sulphuric acid.now gm equivalent of H2SO4 = given mass/equivalent mass = 95.98/49 = 1.95Now until we know the density of the solution we cannot proceed without taking certain assumptions